# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ 3. First examples
In this section, we consider a few general examples of computations of $JT_{\mathbb{G},r,M}(-)$ (some found in [13, §2]). More specific computations for the simpler $JT^{exp}_{\mathbb{G},r,M}(-)$ are given in Section 5. For the examples below, we consider $\mathbb{G}$ to be an infinitesimal group scheme of height $\leq r$ for a specific $r>0$, modifying our notation in these case by using $JT_{\mathbb{G},M}(-)$ to denote $JT_{\mathbb{G},r,M}(-)$.
@@
Let $\mathbb{G}$ be an arbitrary infinitesimal group scheme of height $\leq r$. A $\mathbb{G}$-module $M$ of dimension $m$ is of constant Jordan type $\underline{a}$ if and only if $JT_{\mathbb{G},M}(-)$ is the constant function sending each $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $\underline{a}\in\mathcal{Y}_{\leq,m}$. (See [5] for numerous examples.)
In particular, if $I$ is an injective $\mathbb{G}$-module of dimension $m$, then $p$ divides $m$ and $JT_{\mathbb{G},I}(-)$ sends each $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $\frac{m}{p}[p]\in\mathcal{Y}_{\leq,m}$ (i.e., $JT_{\mathbb{G},I}(\overline{x})=\frac{m}{p}[p]+0[p-1]+\cdots+0[1]$).
@@
Let $\mathbb{G}$ be an arbitrary infinitesimal group scheme of height $\leq r$ and $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$. Consider $0\not=\zeta\in H^{2d}(\mathbb{G},k)$ and take $M$ to be the Carlson $L_{\zeta}$-module as first constructed in [3]; namely, $M$ is the kernel of a map $\tilde{\zeta}:\Omega^{2d}(k)\to k$ representing $\zeta$.
Then $JT_{\mathbb{G},M}(\overline{x})=\frac{n}{p}[p]$ if $((\mu_{x})_{*}\circ\epsilon_{r})^{*}(\zeta))\not=0\in H^{2d}(k(x)[u]/u^{p},k)$, and
$J_{\mathbb{G},M}(\overline{x})=(\frac{n}{p}-1)[p]+[p-1]+[1]$ if $((\mu_{x})_{*}\circ\epsilon_{r})^{*}(\zeta))=0$.
@@
(See [13, Prop 2.9]) Once again, let $\mathbb{G}$ be an arbitrary infinitesimal group scheme of height $\leq r$. Let $M$ be a finite dimensional $\mathbb{G}$-module and $P\twoheadrightarrow M$ a finite dimensional projective cover of $M$ with kernel $\Omega(M)$. If $JT_{\mathbb{G},M}(\overline{x})\ =\underline{a}$ for some $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$, then $JT_{\mathbb{G},\Omega(M)}(\overline{x})\ =\underline{a}^{\perp}+n[p]$ for some integer $n$, where $\underline{a}^{\perp}=\sum_{i=1}^{p-1}a_{p-i}[i]$.
@@
(See [17, §I.3.8].) Consider $\mathbb{G}_{1},\ \mathbb{G}_{2}$, infinitesimal group schemes of height $\leq r$. Let $\mathbb{G}=\mathbb{G}_{1}\times\mathbb{G}_{2}$ and $M_{1}$ be a a finite dimensional $\mathbb{G}_{1}$-module. Set $M$ equal to $ind_{\mathbb{G}_{1}}^{\mathbb{G}}M_{1}\simeq M_{1}\otimes k[\mathbb{G}_{2}]$. Let $\overline{(x,y)}$ be a scheme theoretic point in $\mathbb{P}V_{r}(\mathbb{G})$ represented by the scheme-theoretic point $(x,y)\in V_{r}(\mathbb{G}_{1})\times V_{r}(\mathbb{G}_{2})$. Then
$JT_{\mathbb{G},M}(\overline{(x,0}))\ =\ dim(k[\mathbb{G}_{2}])\cdot JT_{\mathbb{G}_{1},M_{1}}((\overline{x})),$
$JT_{\mathbb{G},M}(\overline{(x,y}))\ =\ dim(M)/p\cdot[p],\quad y\not=0.$
@@
Consider the special case $r=1$. Then applying the functor sending a group scheme to its Lie algebra determines an equivalence between the category of finite group schemes of height 1 and the category of finite dimensional restricted Lie algebra; this sends $\mathbb{G}$ to $\mathfrak{g}$, with $k\mathbb{G}$ identified with the restricted enveloping algebra $\mathfrak{u}(\mathfrak{g})$ of $\mathfrak{g}$. Consequently, a 1-parameter subgroup $\mathbb{G}_{a(1)}\otimes A\to\mathbb{G}_{A}$ corresponds to a choice of element $\alpha\in A\otimes\mathfrak{g}$ with $\alpha^{[p]}=0$. Thus, $V_{1}(\mathbb{G})$ is identified with the $p$-nilpotent cone $\mathcal{N}_{p}(\mathfrak{g})\subset\mathfrak{g}$.
Let $M$ be a faithful representation of $\mathbb{G}$ of dimension $N$, and associate to $M$ an embedding $\phi:\mathbb{G}\hookrightarrow GL_{N(1)}\hookrightarrow GL_{N}$ with the property that $M=\phi^{*}(V_{N})$ (where $V_{N}$ is the defining representation for $GL_{N}$). For any $\alpha\in\mathcal{N}_{p}(\mathfrak{g})(A)$, the associated 1-parameter subgroup of $\mathbb{G}$ sends a $p$-nilpotent element $s\in A$ to the $N\times N$ matrix $1+s\cdot\alpha+\frac{s^{2}}{2}\alpha^{2}+\cdots+\frac{s^{p-1}}{(p-1)!}\alpha^{p-1}\in GL_{N}(A)$ (which is in the image of $\mathbb{G}(A)\to GL_{N}(A)$). Moreover, if the scheme-theoretic point $x\in\mathcal{N}_{p}(\mathfrak{g})$ corresponds to $\alpha\in k(x)\otimes\mathfrak{g}$, then $JT_{\mathbb{G},M}(x)$ is the Jordan type of $\alpha\in\mathcal{N}_{p}(\mathfrak{g})(A)\subset M_{N,N}(A)$.
In particular, if $M=\mathfrak{g}l_{N}$, the adjoint representation of $GL_{N}$, then $JT_{GL_{N},1,\mathfrak{g}l_{N}}(-)$ is the class function taking a $p$-nilpotent element of $\mathfrak{g}l_{N}$ to its Jordan block structure viewed as a Young diagram.
@@
Assume $\mathbb{G}=\mathbb{G}_{a(r)}$ for some $r>0$. Then $\mathbb{G}$ admits an embedding in $GL_{2}$ (as the $r$-th Frobenius kernel of the unipotent subgroup $U_{2}$ of strictly upper triangular matrices). As in Example 2.6, $\mathbb{P}V_{r}(\mathbb{G})$ is a weighted projective space with weights $(1,p,\cdots,p^{r-1})$ associated to the graded polynomial algebra $k[T_{0},T_{1},\ldots,T_{r-1}]$ where $T_{i}$ has weight $p^{i}$. A scheme-theoretic point $\underline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ lifts to a scheme-theoretic point $x\in V_{r}(X)$ represented by a non-zero $r$-tuple $\underline{a}=(a_{0},\ldots,a_{r-1})$ of elements in $k(x)$ with associated 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{a(r),k(x)}\simeq k(x)[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$. An explicit (but quite complicated) formula for $\mu_{x*}(u_{r-1})$ for $x$ represented by $(a_{0},\ldots,a_{r-1})$ is given in [23, (6.5.12)].
The sum of terms linear in $u_{i}$ in the expression for $\mu_{x*}(u_{r-1})$ is $\sum_{i=0}^{r-1}a_{r-1-i}^{p^{i}}u_{i}\in k(x)\otimes k\mathbb{G}$. See Example 2.10.
@@ 4. Exponential Jordan type
In this section, we consider affine group schemes whose 1-parameter subgroups are given by exponentiating $r$-tuples of $p$-nilpotent, pair-wise commuting elements of $\mathfrak{g}\equiv Lie(\mathbb{G})$. The more explicit nature of these 1-parameter subgroups allow us to simplify our Jordan type functors, not necessarily preserving their values but preserving their close relationship with support varieties. The key to this simplification is a fundamental computation of [15, Thm 1.12] which tells us that only linear terms of commuting generators of group actions given as polynomials in these generators are needed to determine support varieties.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. The functor on commutative $k$ algebras sending $A$ to an $r$-tuple $\underline{B}=(B_{0},\ldots,B_{r-1})$ of elements in $A\otimes\mathfrak{g}$ which satisfy the conditions that $B_{s}^{[p]}=0$ for all $i,0\leq s<r$ and that $[B_{s},B_{s^{\prime}}]=0$ for all $s,s^{\prime}$ with $0\leq s,s^{\prime}<r$ is represented by an affine scheme $\mathcal{C}_{r}(\mathfrak{g})$. In contrast to [11, 1.11.2], we give $\mathcal{C}_{r}(\mathfrak{g})$ the monoid action $\mathbb{A}^{1}\times\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{C}_{r}(\mathfrak{g})$ by sending $(\alpha;B_{0},\ldots,B_{r-1})$ to $(\alpha\cdot B_{0},\alpha^{p}\cdot B_{1},\ldots,\alpha^{p^{r-1}}\cdot B_{r-1})$.
For $\mathbb{G},\mathcal{E}_{(r)}$ an affine group scheme of exponential type of height $r$ as in Definitions 4.2 and 4.6, the equality $(\mathcal{U}_{\mathfrak{g},r})_{*}(u_{r-1})\ =(\Phi_{\mathfrak{g},r}\otimes 1)^{*}(\Theta_{\mathbb{G},r})$ of Proposition 4.5(2) below gives an identification of the universal $p$-nilpotent operator $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes\mathbb{G}_{(r)}$ utilizing $\mathcal{U}_{\mathfrak{g},r}$ of (4.3.2) in place of the universal height $r$ 1-parameter subgroup $\mathcal{U}_{\mathbb{G},r}$ This identification has the advantage that $\mathcal{U}_{\mathfrak{g},r}$ is given in terms of the “explicit” representation of 1-parameter subgroups of $\mathbb{G}_{(r)}$ in terms of $r$-tuples of $p$-nilpotent, pairwise elements of $\mathfrak{g}$.
Definition 4.2.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. Fix a positive integer $r$. Assume that $\mathbb{G}$ is equipped with a morphism
@@
which satisfies the following conditions for any commutative $k$-algebra $A$, any $p$-nilpotent $B\in A\otimes\mathfrak{g}$, and any $\alpha,\beta\in\mathbb{G}_{a(r)}(A)$:
@@
    The specialization at $B$, $\mathcal{E}_{(r),B}:\mathbb{G}_{a(r),A}\to\mathbb{G}_{(r),A}$, is a 1-parameter subgroup of $\mathbb{G}_{(r),A}$.
@@
    If $B^{\prime}\in A\otimes\mathfrak{g}$ commutes with $B$, then
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ equipped with $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$ satisfying the conditions of Definition 4.2. For any $s,0\leq s<r$, we define
(4.3.1) $\mathcal{E}_{r|s}\quad\equiv\quad\mathcal{E}_{(r),B}\circ(1\times p_{s})\circ(F^{s}\times 1)$
to be the composition
$\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g})\ \to\ \mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g})\to\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$
where $p_{s}$ is projection onto the $s$-th factor and $F^{s}$ is the $s$-th iterate of the Frobenius morphism.
@@
$\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g})\ \to(\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g}))^{\times r}\ \to\ (\mathbb{G}_{(r)})^{\times r}\ \to\mathbb{G}_{(r)}$

### v2

@@ 3. First examples
In this section, we consider a few general examples of computations of $JT_{\mathbb{G},r,M}(-)$ (some found in [15, §2]). More specific computations for the simpler $JT^{exp}_{\mathbb{G},r,M}(-)$ are given in Section 5.
@@
Let $\mathbb{G}=\mathbb{G}_{(r)}$ be an infinitesimal group scheme of height $\leq r$. A $\mathbb{G}$-module $M$ of dimension $m$ is of constant Jordan type $\underline{a}$ if and only if $JT_{\mathbb{G},M,r}(-)$ is the constant function sending each $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $\underline{a}\in\mathcal{Y}_{m}$. (See [7] for numerous examples.)
In particular, if $I$ is an injective $\mathbb{G}$-module of dimension $m$, then $p$ divides $m$ and $JT_{\mathbb{G},I,r}(-)$ sends each $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $\frac{m}{p}[p]\in\mathcal{Y}_{m}$ (i.e., $JT_{\mathbb{G},I-r}(\overline{x})=\frac{m}{p}[p]+0[p-1]+\cdots+0[1]$).
@@
Let $\mathbb{G}=\mathbb{G}_{(r)}$ be an infinitesimal group scheme of height $\leq r$ and $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$. Consider $0\not=\zeta\in H^{2d}(\mathbb{G},k)$ and take $M$ to be the Carlson $L_{\zeta}$-module as first constructed in [5]; namely, $M$ is the kernel of a map $\tilde{\zeta}:\Omega^{2d}(k)\to k$ representing $\zeta$.
Then the dimension $n$ of $M$ is divisible by $p$. Moreover,
$J_{\mathbb{G},M,r}(\overline{x})=\frac{n}{p}[p]\quad{\text{i}f}\quad((\mu_{x})_{*}\circ\epsilon_{r})^{*}(\zeta))\not=0\in H^{2d}(k(x)[u]/u^{p},k),$
$J_{\mathbb{G},M,r}(\overline{x})=(\frac{n}{p}-1)[p]+[p-1]+[1]\quad{\text{i}f}\quad((\mu_{x})_{*}\circ\epsilon_{r})^{*}(\zeta))=0\in H^{2d}(k(x)[u]/u^{p},k).$
@@
(See [15, Prop 2.9]) Let $\mathbb{G}=\mathbb{G}_{(r)}$ be an infinitesimal group scheme of height $\leq r$. Let $M$ be a finite dimensional $\mathbb{G}$-module and $P\twoheadrightarrow M$ a finite dimensional projective cover of $M$ with kernel $\Omega(M)$. If $JT_{\mathbb{G},M,r}(\overline{x})\ =\underline{a}$ for some $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$, then $JT_{\mathbb{G},\Omega(M),r}(\overline{x})\ =\underline{a}^{\perp}+n[p]$ for some integer $n$, where $\underline{a}^{\perp}=\sum_{i=1}^{p-1}a_{p-i}[i]$.
@@
(See [20, §I.3.8].) Consider $\mathbb{G}_{1},\ \mathbb{G}_{2}$, infinitesimal group schemes of height $\leq r$. Let $\mathbb{G}=\mathbb{G}_{1}\times\mathbb{G}_{2}$ and $M_{1}$ be a a finite dimensional $\mathbb{G}_{1}$-module. Set $M$ equal to $ind_{\mathbb{G}_{1}}^{\mathbb{G}}M_{1}\simeq M_{1}\otimes k[\mathbb{G}_{2}]$. Let $\overline{(x,y)}$ be a scheme-theoretic point in $\mathbb{P}V_{r}(\mathbb{G})$ represented by the scheme-theoretic point $(x,y)\in V_{r}(\mathbb{G}_{1})\times V_{r}(\mathbb{G}_{2})$. Then
$JT_{\mathbb{G},M,r}(\overline{(x,0}))\ =\ dim(k[\mathbb{G}_{2}])\cdot JT_{\mathbb{G}_{1},M_{1}}((\overline{x})),$
$JT_{\mathbb{G},M,r}(\overline{(x,y}))\ =\ dim(M)/p\cdot[p],\quad y\not=0.$
We generalize the previous example as follows.
@@
Consider $\mathbb{G}_{1},\ \mathbb{G}_{2}$, infinitesimal group schemes of height $\leq r$, a finite dimensional $\mathbb{G}_{1}$-module $M_{1}$, and a finite dimensional $\mathbb{G}_{2}$-module $M_{2}$. Let $(x,y)$ be a scheme-theoretic point in $V_{r}(\mathbb{G}_{1})\times V_{r}(\mathbb{G}_{2})$. Then
$JT_{\mathbb{G}_{1}\times\mathbb{G}_{2},M\otimes N,r}((x,y))\quad=\quad JT_{\mathbb{G}_{1},M,r}(x)\otimes JT_{\mathbb{G}_{2},N,r}(y).$
@@
Let $\mathbb{G}$ be an infinitesimal group scheme of height 1 with Lie algebra $\mathfrak{g}$. Thus, $k\mathbb{G}$ is naturally identified with $\mathfrak{u}(\mathfrak{g})$, the restricted enveloping algebra of $\mathfrak{g}$ and $V_{1}(\mathbb{G})$ is naturally identified with the $p$-nilpotent cone $\mathcal{N}_{p}(\mathfrak{g})\subset\mathfrak{g}$. (See [26, Lem 1.6].)
Let $V$ be a faithful representation of $\mathbb{G}$ of dimension $N$, and associate to $V$ a closed embedding $\phi:\mathbb{G}\hookrightarrow GL_{N(1)}\hookrightarrow GL_{N}$. For any $\alpha\in\mathcal{N}_{p}(\mathfrak{g})(A)$, the associated 1-parameter subgroup of $\mu_{\alpha}:\mathbb{G}_{a(1),A}\to\mathbb{G}_{A}$ sends $s\in\mathbb{G}_{a(1)}(A)$ (i.e., $s$ a $p$-nilpotent element of $A$) to
$\mu_{\alpha}(s)\ =1+s\cdot\alpha+\frac{s^{2}}{2}\alpha^{2}+\cdots+\frac{s^{p-1}}{(p-1)!}\alpha^{p-1}\ \in\ \mathbb{G}_{(1)}(A)\ \subset\ GL_{N}(A),$
where the powers of $\alpha$ are taken in $\mathfrak{g}l_{N}$. We associate to $\mu_{\alpha}$ the linear operator $(\mu_{\alpha})_{*}:A\otimes k\mathbb{G}_{a(1)}\to\mathcal{A}\otimes k\mathbb{G}$.
Let $M$ be a finite dimensional $\mathbb{G}$-module of dimension $m$ and choose a scheme-theoretic point $\alpha\in\mathcal{N}_{p}(\mathfrak{g})$, given by $\alpha\in k(\alpha)\otimes\mathcal{N}_{p}(\mathfrak{g})$. Then $JT_{\mathbb{G},M,1}(\alpha)$ is the Jordan type of $(\mu_{\alpha})_{*}(1\otimes u)\in k(\alpha)\otimes k\mathbb{G}$ acting on $k(\alpha)\otimes M$. This action is the action of $\alpha$ as an element of the Lie algebra $k(\alpha)\otimes\mathfrak{g}$ action on $k(\alpha)\otimes M$.
Example 3.7.
Assume $\mathbb{G}=\mathbb{G}_{a(r)}$ for some $r>0$. Then $\mathbb{G}$ admits an embedding in $GL_{2}$ as the $r$-th Frobenius kernel of the unipotent subgroup $U_{2}$ of strictly upper triangular matrices. As in Example 2.10, $\mathbb{P}V_{r}(\mathbb{G})$ is a weighted projective space with weights $(1,p,\cdots,p^{r-1})$ associated to the graded polynomial algebra $k[T_{0},T_{1},\ldots,T_{r-1}]$ where $T_{i}$ has weight $p^{i}$. Consider a non-zero $r$-tuple $\underline{a}=(a_{0},\ldots,a_{r-1})$ of elements of $k(x)$ determining a scheme-theoretic point $x\in V_{r}(X)$ projecting to the scheme-theoretic point $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$, and consider the associated 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{a(r),k(x)}.$ An explicit (but quite complicated) formula for $\mu_{x*}(1\otimes u_{r-1})\in k(x)\otimes k\mathbb{G}_{a(r)}\ =\ k(x)\otimes k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$ is given in [27, (6.5.12)].
The sum of terms linear in the $u_{i}$’s in the expression for $\mu_{x*}(1\otimes u_{r-1})$ is $\sum_{i=0}^{r-1}a_{r-1-i}^{p^{i}}u_{i}\in k(x)\otimes k\mathbb{G}$. Replacing $\mu_{x*}(1\otimes u_{r-1})$ by $\sum_{i=0}^{r-1}a_{r-1-i}^{p^{i}}u_{i}$ is a special case of replacing $\Theta_{\mathbb{G},r}$ by $\Theta^{exp}_{\mathfrak{g},r}$ as discussed in Section 4.
@@ 4. Exponential Jordan type
In this section, we consider affine group schemes whose 1-parameter subgroups are given by exponentiating $r$-tuples of $p$-nilpotent, pair-wise commuting elements of $\mathfrak{g}\equiv Lie(\mathbb{G})$. The more explicit nature of these 1-parameter subgroups allows us to simplify our Jordan type functions, not necessarily preserving their values but preserving their close relationship with support varieties.
For an affine group scheme $\mathbb{G}$ equipped with an exponential map $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$ of height $r$, there is a natural morphism $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ V_{r}(\mathbb{G})$ where $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is the affine variety of $r$-tuples of $p$-nilpotent, pair-wise commuting elements of $\mathfrak{g}$ made explicit in Definition 4.2. We say that $\mathbb{G}$ is of exponential type of height $r$ if $\Phi_{\mathfrak{g},r}$ is a homeomorphism of scheme-theoretic points. Examples include Frobenius kernels of various classical groups, their parabolic subgroups, and unipotent subgroups of these parabolic subgroups. (See, for example, [25].) For $\mathbb{G}$ of exponential type of height $r$, Theorem 4.8 introduces $\Theta_{\mathfrak{g},r}\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}$ which very closely models $\Theta_{\mathbb{G},r}$.
This leads to our simpler $p$-nilpotent operator $\Theta_{\mathfrak{g},r}^{exp}\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\ \otimes k\mathbb{G}_{(r)}$ for $\mathbb{G}$ of exponential type of height $r$, a linearization of $\Theta_{g,r}$. A fundamental computation of [17, Thm 1.12] enables us to conclude that $\Theta_{\mathfrak{g},r}$ and $\Theta_{\mathfrak{g},r}^{exp}$ determine the same support varieties for finite dimensional $\mathbb{G}_{(r)}$-modules. By adopting the formalism we have discussed for $JT_{\mathbb{G},r,M}(-)$, we establish the Jordan type functions $JT_{\mathfrak{g},r,M}(-)$ and $JT_{\mathfrak{g},r,M}^{exp}(-)$ which satisfy similar properties.
We begin by defining the class of affine group schemes $\mathbb{G}$ for which $r$-tuples of $p$-nilpotent, pair-wise commuting elements of the Lie algebra of $\mathbb{G}$ determine height $r$ 1-parameter subgroups of $\mathbb{G}$.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. Fix a positive integer $r$. We say that a morphism
@@
equips $\mathbb{G}$ with the structure of an exponentiation of height $r$ if $\mathcal{E}_{(r)}$ satisfies the following conditions for any commutative $k$-algebra AA, $p$-nilpotent elements $B,B^{\prime}\ \in A\otimes\mathfrak{g}$, and elements $\alpha,\beta\in\mathbb{G}_{a(r)}(A)$:
@@
    The specialization of $\mathcal{E}_{(r)}$ along $B:k[\mathcal{N}_{p}(\mathfrak{g})]\to A$, $\mathcal{E}_{(r),B}:\mathbb{G}_{a(r),A}\to\mathbb{G}_{(r),A}$, is a 1-parameter subgroup of $\mathbb{G}_{(r),A}$.
@@
    If $B^{\prime}$ commutes with $B$, then
@@
As exhibited in Construction 4.3, the structure $\mathcal{E}_{(r)}$ of exponentiation of height $r$ for $\mathbb{G}$ enables us to “exponentiate” $r$-tuples of $p$-nilpotent, pair-wise commuting elements of $\mathfrak{g}$. We first define the variety of such $r$-tuples, $\mathcal{C}_{r}(\mathfrak{g})$.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. Consider the functor on commutative $k$ algebras sending $A$ to an $r$-tuple $\underline{B}=(B_{0},\ldots,B_{r-1})$ of elements in $A\otimes\mathfrak{g}$ which satisfies the conditions that $B_{s}^{[p]}=0$ for all $s,\ 0\leq s<r$ and that $B_{s},B_{s^{\prime}}]=0$ for all $s,s^{\prime}$ with $0\leq s,s^{\prime}<r$. This functor is represented by an affine scheme $\mathcal{C}_{r}(\mathfrak{g})$. In contrast to [13, 1.11.2], we give $\mathcal{C}_{r}(\mathfrak{g})$ the monoid action $\mathbb{A}^{1}\times\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{C}_{r}(\mathfrak{g})$ by sending $(\alpha;B_{0},\ldots,B_{r-1})$ to $(\alpha\cdot B_{0},\alpha^{p}\cdot B_{1},\ldots,\alpha^{p^{r-1}}\cdot B_{r-1})$.
We now assume the $\mathbb{G}$ is equipped with an exponential $\mathcal{E}_{(r)}$ of height $r$. The following construction of $\mathcal{E}_{\mathfrak{g},r}:\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathbb{G}_{(r)}$ given in (4.3.2) extends the exponential structure $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\ \to\ \mathbb{G}_{(r)}$. This leads to the “Lie algebra analogue”
$\mathcal{U}_{\mathfrak{g},r}:\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad\mathbb{G}_{(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$
of the universal 1-parameter subgroup $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r)}\times V_{r}(\mathbb{G})\to\mathbb{G}_{(r)}\times V_{r}(\mathbb{G})$ of Definition 2.12.
Construction 4.3.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ equipped with an exponentiation $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$ of height $r$. For any $s,\ 0\leq s<r$, we define
(4.3.1) $\mathcal{E}_{r|s}\quad\equiv\quad\mathcal{E}_{(r)}\circ(1\times pr_{s})\circ(F^{s}\times 1):\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g})\ \to\ \mathbb{G}_{(r)}$
where $pr_{s}$ is projection onto the $s$-th factor and $F^{s}$ is the $s$-th iterate of the Frobenius morphism. We denote by $\mathcal{E}_{r|s,B}$ the restriction of $\mathcal{E}_{r|s}$ along a (geometric) point $B:k(B)\to\mathcal{N}_{p}(\mathfrak{g})$.
@@
$\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to(\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})))^{\times r}\ \to\ (\mathbb{G}_{(r)})^{\times r}\ \to\mathbb{G}_{(r)}$